OpenStudy (anonymous):

Convert each degree measure into radians. 50° -300° 670° 220°

OpenStudy (anonymous):

Here\[radians = degree \times \frac{ \pi }{ 180 }\]

OpenStudy (anonymous):

so how would i figure that out im still confused.

OpenStudy (anonymous):

Your first one for example would be \[radians = 50 \times \frac{ \pi }{ 180 }\] Just put into a calculator

OpenStudy (anonymous):

That will cancel down to \[\frac{ 5 \pi}{ 18 } radians\]

OpenStudy (anonymous):

but i dont have a calculator what about the next one.

OpenStudy (anonymous):

same thing. Sorry i don't really know an easier way to do this. but the next one would be \[-\frac{ 5 \pi }{ 3 }\] It's just a matter of cancelling down fractions really as you can leave pi in the answer for radians.

OpenStudy (anonymous):

pi radians = 180 degrees, from this you can sort of guess and trial and error